# The StudyChat Dataset: Student Dialogues With ChatGPT in an Artificial Intelligence Course

### v2

@@ Abstract
The widespread availability of large language models (LLMs), such as ChatGPT, has significantly impacted education, raising both opportunities and challenges. Students can frequently interact with LLM-powered, interactive learning tools, but their usage patterns need to be analyzed to ensure ethical usage of these tools. To better understand how students interact with LLMs in an academic setting, we introduce StudyChat, a publicly available dataset capturing real-world student interactions with an LLM-powered tutoring chatbot in a semester-long, university-level artificial intelligence (AI) course. We deploy a web application that replicates ChatGPT’s core functionalities, and use it to log student interactions with the LLM while working on programming assignments. We collect 1,197 conversations, which we annotate using a dialogue act labeling schema inspired by observed interaction patterns and prior research. Additionally, we analyze these interactions, highlight behavioral trends, and analyze how specific usage patterns relate to course outcomes. StudyChat provides a rich resource for the learning sciences and AI in education communities, enabling further research into the evolving role of LLMs in education  [^1] https://huggingface.co/datasets/wmcnicho/StudyChat.
@@ 1 Introduction
Advances in large language models (LLMs) have had major ramifications on almost all levels of education. Many LLM-powered educational technologies have recently been developed, leveraging the generative capabilities of these models to create new course content (MacNeil et al., 2022), assignments (MacNeil et al., 2023), and interactive learning experiences(Jin et al., 2024; Taneja et al., 2024). These interactive technologies enable rich engagement with course material by allowing students to engage in on-demand conversations with an interactive tutor, and receive personalized feedback on their work. However, while the potential use cases for LLMs in education are vast, many educators are understandably concerned about student misuse of LLM chatbot tools, such as ChatGPTLau and Guo (2023). While a motivated learner can use these tools in a learning positive way, an unmotivated student could use them to complete their assignments on their behalf, circumventing the learning objectives of the assignment, or become over-reliant on the model(Bastani et al., 2024).
@@
This dichotomy between learning opportunity and academic risk necessitates an exploration into understanding student behavior with LLM chatbot tools. A more nuanced understanding of this behavior could inform further development of LLM-based Intelligent Tutoring Systems (ITS) that dynamically adapt to students, encouraging learning-focused behaviors and discouraging counterproductive ones. Furthermore, a deeper understanding of student behavior can lead to more robust student models, which better estimate knowledge states, misconceptions, and learning trajectories. For example, such a model could potentially detect when a student is overly reliant on LLM responses and flag this to instructors in real-time, or be used to help develop assessments which are specifically designed to test for this over-reliance.
@@
In this work, we conduct a case study to explore student behavior with an LLM, specifically in the context of upper-division, university-level Computer Science (CS) education. We develop a web application with the same core features and underlying LLM as ChatGPT, and deploy it in an introductory-level course on artificial intelligence (AI). We encourage students to use this tool, without restrictions, on programming assignments over the course of the semester. With student consent, we recorded over 1,000 student-LLM conversations from 84 students across 7 unique assignments, completed over 15 weeks. We also collect 455 assignment submissions from 65 students who further agreed to share their work. In addition, we annotate these conversation turns with a dialogue act labeling schema; the schema is inspired by other existing studies on co-programming behavior and our observations during a systematic review of the conversations. We leverage an LLM prompting-based approach to perform this annotation at scale and validate this approach with a human evaluation.
To explore how students engage with an LLM during programming assignments, we analyze these dialogue act labels and highlight notable trends in behavior. In addition, we investigate if certain patterns of behavior correlate with course performance by performing a regression analysis. Our key findings are 1.) students primarily use the LLM to ask questions, such as assignment clarification or code explanation, but co-programming and answer seeking behaviors are also very common 2.) certain engagement patterns such as asking conceptual questions or actively discussing problem-solving approaches with the LLM, correlate with better scores on graded items, and 3.) students who tend to use the LLM for debugging and tool assistance tend to perform worse on tool-prohibited items which perhaps highlights LLM over-reliance or under-preparedness. These findings provide data-driven insights into how students interact with LLMs. We release this to the broader research community to foster collaboration and exploration into LLM’s impact on CS education.
@@ 2.1 ChatGPT Classroom Studies
Closely related to this work are a handful of cases studies exploring student use of ChatGPT in university level CS classrooms. Ghimire et. al explore student use of ChatGPT for two assignments in a CS1 programming course, finding students often ask questions of the system instead of directly seeking solutions (Ghimire and Edwards, 2024). Ma et. al also perform a case study where CS1 students were allowed to use ChatGPT freely on in-class exercises and on a course project, but only for conceptual questions on homework(Ma et al., 2024). Additionally, Brender et. al perform a case study on how graduate students in a robotics course utilize ChatGPT on a voluntary assignment, controlling for learning outcomes and clustering behavior into broad categories (Brender et al., 2024). Analysis of student prompts in these studies have noted a student trend to use models to ask questions more than directly writing code solutions which is consistent with our observations and those by Bhalerao (2024).Surveys on student perceptions towards LLMs have also been conducted which suggest that CS student adoption of LLMs is nuanced and not universal(Weber et al., 2024). Similarly, we observe that a some students use the system heavily while most students engage moderately with the model, revealing a range of adoption behaviors.
@@ 2.2 Education Discourse Analysis
It has been demonstrated by many that tutoring is a highly effective form of instruction Cohen et al. (1982). Thus, human tutoring dialogue has been studied extensively Graesser et al. (1995). In recent years, dialogue act (DA) classification on conversation turns is often used in this discourse analysisChen et al. (2011). DA labels classify a participant’s utterance by the function it serves in the conversation. When DA labeling is applied to an entire corpus of dialogues, it enables discourse-level analysis into patterns of conversation that are most frequent or significant. In an education context, significant patterns are often those associated with better learning outcomes, as such existing works have studied such patterns in tutor-student or student-student learning environmentsGraesser et al. (1995); Earle-Randell et al. (2023). Existing works also analyze human-machine classroom conversations between students and ITS chatbots Vail and Boyer (2014); Graesser et al. (2004); Fisher et al. (2020). These works provide promising insights which can inform more effective tutoring pedagogyRus et al. (2017) and interactive learning toolsChen et al. (2011).
@@
While DA analysis is an attractive approach to discourse analysis, the labeling process poses many challenges, solutions to which constitute an large body of workBoyer et al. (2011); Tan et al. (2023); Ezen-Can and Boyer (2015); Lin et al. (2023); Boyer et al. (2010); Kumaran et al. (2023). First, there exists the task of annotating a large corpus of conversations. To automate such a task, various machine learning based approaches to this have been used such as Hidden Markov ModelsBoyer et al. (2010), decision treesSamei et al. (2014), and Markov Random FieldsEzen-Can and Boyer (2015). In our work, we employ a LLM prompt-based approach to dialogue act labeling, similar to that ofWang et al. (2023). Second, there exists the challenge of robust schema design, where existing work has looked into refining prior labeling schemata to be less ambiguousVail and Boyer (2014) and to try to develop classroom agnostic labeling schemaHennessy et al. (2016).
2.3 Education Dialog Datasets
@@
In an effort to collectively improve data-driven learning technologies, there have been recent efforts to expand the amount of education dialog datasets in a variety of learning domains. Datasets such as TalkMoves and NCTE directly capture and annotate classroom discourse in K-12 lessons Suresh et al. (2022); Demszky and Hill (2022). Other works employ crowd workers or leverage LLMs to simulate the student or teacher role, such as MathDialMacina et al. (2023), Teacher-Student Chatroom Corpus (TSCC)Caines et al. (2020), Conversational Instruction with Multi-responses and Actions (CIMA)Stasaski et al. (2020), and GrounDialogZhang et al. (2023). Our work contributes to this growing body of research by providing a large-scale, annotated dialog dataset of in a university-level CS course, offering insights into how students engage with LLM-chatbot tools in a real learning environment.
@@ 3 Dataset
In this section, we detail the process by which the StudyChat dataset was formed via the classroom context, application, and dialogue act annotation process.
@@
Conceptual Questions Programming Language, Python Library, Computer Science, Programming Tools, Other Concept The user asks a general question about programming, libraries, tools, or CS concepts unrelated to the conversation.
@@ 3.1 Participants and Course Context
Our IRB-approved study was conducted among students aged 18 or above, who were enrolled in an AI course at a US R1 university in the Fall 2024 semester. 117 students were enrolled in the course, of which 84 agreed to participate in the study. Students who did not participate were also given equivalent access to use the tool over the semester, but their data was not collected. Each week, students attend two 75-minute lectures and independently complete assignments related to the topics covered in lecture. The course consists of 3 modules taught over 15 weeks. During each module, students complete 2-3 Python programming assignments related to the lecture material. In addition to the assignments, students were assessed via 2 end-of-module exams, a final exam, and a group project.
@@ 3.2 Data Collection
In order to effectively collect student interactions, we developed and deployed a web application designed to function similar to consumer LLM chatbot interfaces, as shown in Figure 1. As our backing chatbot LLM, we used gpt-4o-mini, the LLM for the free tier of ChatGPT at the time of this studygpt . Students were instructed they could use this ChatGPT-based tool, without restrictions, on all assignments in the course and the final project. It was emphasized that their approach to using the tool would not affect their grade and they were encouraged to use in any manner of their choosing. We used the default "You are a helpful assistant" system prompt, which does not discouraging answer-seeking behavior, and did not provide the chatbot with any course materials for reference. The reasoning for these decisions was to encourage students to use our platform instead of seeking an outside application for homework help and that the collected conversations would more accurately reflect how students generally use non-instructional, consumer chatbots. In addition to the conversations, we also collected the 455 submissions by 65 of the 84 student participants consented to share their assignment solutions. Finally we collected responses to a pre-study survey on students’ perspectives on LLM-based programming tools.
@@ 3.3 Filtering of Student Identifiable Information
To remove any Personally Identifiable Information (PII) from the conversations, we systematically analyze the conversations to uncover patterns of PII exposure. We first randomly sample 12 conversations from each course topic and read through the conversations, scanning for traces of PII. Our main finding of potentially sensitive student information is in the form of pasted terminal outputs and error messages, which sometimes contain PII, such as a user’s directory or GitHub account name. Compiling these common patterns, we then develop a regular expression-based script to remove these patterns and any direct matches to known user PII in the course roster. Across the entire dataset, this process removes 2009 cases of possible PII exposure. Then, we sample 12 additional conversations from each topic and performed the same procedure. In our second pass of the second set of conversations, we did not find any additional patterns or cases of PII exposure. We acknowledge that although seemingly effective, this process does not guarantee the removal of all PII exposures.
Figure 2: Violin plot showing the distribution of total utterances per user for each assignment topic. The width of each violin represents the density of user interactions, while inner quartiles indicate the spread of engagement.
@@ 3.4 Data Statistics
Across all conversations, we collect 7,856 student utterances and corresponding LLM responses over 7 graded assignments. The mean conversation length in the dataset is 7.07 utterances, with a median of 3 utterances per conversation session. On average, each student has 23.13 utterances and 13.5 distinct conversations with the chatbot. Figure 2 shows the number of utterances for each user across topics, where there is an average of 70.89 utterances for each topic. We see a skewed distribution where a small group of students use the model extensively, over 100 utterances for some assignments, while the average median number of utterances per student is 7.7. Specifically, we find that the top 5% most-active students account for 21.82% of all utterances in the dataset.
@@ 3.5 Dialogue Act Annotation
In order to better understand common behavior and discourse-level trends in the conversations, we develop a dialogue act labeling schema for student turns of the conversations. The schema, as shown in Table 1 is two-leveled and captures common user behaviors in the dataset. This scheme is developed on top of existing work Wood et al. (2018) which conducted a “Wizard of Oz” experiment with programmers who were debugging alongside a simulated virtual assistant. It also mirrors our own observations after systematically reviewing the conversations, and aligns with those of Ma. et alMa et al. (2024).
@@
First, we sample a set of 150 utterances and develop an initial schema with labeling guidelines, including examples and descriptions for each dialogue act. We then provide them to a research assistant, who independently labels the same set given our initial schema. We iterate on our labeling instructions and schema by computing the agreement between the two annotators and identifying sources of disagreement. After this iteration, our inter-rater agreement reached a Cohen’s kappaCohen (1960) of 0.910 for broad level labels and 0.788 for specific labels, indicating almost perfect agreement on the broad category of dialogue act and substantial agreement on the specific label Landis and Koch (1977). To validate this labeling, we provided these final instructions to two additional annotators who independently label the same set of utterances. We observed and average 4-way inter-rater agreement Cohen’s kappa of 0.740 for broad level agreement and 0.575 for specific level agreement, indicating substantial agreement on broad level labeling and moderate agreement on specific labeling.
@@
To apply this labeling at scale to the reminder of the dataset and since manual labeling is extraneous, we perform an LLM-prompting based approach similar to prior workWang et al. (2023). We prompt the LLM with detailed instructions on the task, the annotation instructions, and the entire conversation up until the student turn to be labeled, and ask the LLM to provide the label based on our schema, along with a justification. We utilize GPT4o to perform this labeling, which we applied to the entire PII-filtered dataset.
@@
To validate the LLM labels, we again systematically sample 154 utterances, distinct from those used in developing the initial labeling schema. We provide them to the same 4 human annotators without any further discussion or instructions. The resulting human inter-annotator agreement on the held out set is 0.578 on broad labels and 0.538 on specific labels, indicating moderate agreement between human annotators on the held out set. This drop in agreement is expected, since the new utterances in the held out set present new ambiguous utterances to which multiple labels may apply. The average human-LLM agreement on the held out set is 0.446 on the broad labels and 0.333 on the specific labels, indicating moderate and fair agreement respectively. There are many potential causes for the disparity between human and LLM labeling; the most likely reason is that the LLM does not have knowledge of the course context like the human annotators, who have either taught the course or completed it as a student in prior semesters.
Overall, we emphasize that there are challenges in creating a universal dialogue act labeling schema for any domain, and acknowledge that there is room for further refinement, which we leave for future work since it is not the focus of the current paper, and is the subject of many related works Boyer et al. (2011); Tan et al. (2023); Ezen-Can and Boyer (2015); Lin et al. (2023); Boyer et al. (2010); Kumaran et al. (2023).

### v3

# The StudyChat Dataset: Exploring Student Dialogues With ChatGPT in an Artificial Intelligence Course

@@ Abstract
The widespread availability of large language models (LLMs), such as ChatGPT, has significantly impacted education, raising both opportunities and challenges. Students can frequently interact with LLM-powered, interactive learning tools, but their usage patterns need to be monitored and understood. We introduce StudyChat, a publicly available dataset capturing real-world student interactions with an LLM-powered tutoring chatbot in a semester-long, university-level artificial intelligence (AI) course. We deploy a web application that replicates ChatGPT’s core functionalities, and use it to log student interactions with the LLM while working on programming assignments. We collect 16,851 interactions, which we annotate using a dialogue act labeling schema inspired by observed interaction patterns and prior research. We analyze these interactions, highlight usage trends, and analyze how specific student behavior correlates with their course outcome. We find that students who prompt LLMs for conceptual understanding and coding help tend to perform better on assignments and exams. Moreover, students who use LLMs to write reports and circumvent assignment learning objectives have lower outcomes on exams than others. StudyChat serves as a shared resource to facilitate further research on the evolving role of LLMs in education.  [^1] https://huggingface.co/datasets/wmcnicho/StudyChat.
@@ 1 Introduction
Advances in large language models (LLMs) have had major ramifications on almost all levels of education. Many LLM-powered educational technologies have recently been developed, leveraging the generative capabilities of these models to create new course content (MacNeil et al., 2022), assignments (MacNeil et al., 2023), and interactive learning experiences (Jin et al., 2024; Taneja et al., 2024). These interactive technologies enable rich engagement with course material by allowing students to engage in on-demand conversations with an interactive tutor, and receive personalized feedback on their work. However, while the potential use cases for LLMs in education are vast, many educators are understandably concerned about student misuse of LLM chatbot tools, such as ChatGPT Lau and Guo (2023). While a motivated learner can use these tools positively, an unmotivated student could use them to complete their assignments, circumventing the learning objectives of the assignment, or become over-reliant on the model (Bastani et al., 2024).
@@
This dichotomy between learning opportunity and academic risk necessitates an exploration into understanding student behavior with LLM chatbot tools. A more nuanced understanding of this behavior could inform further development of LLM-based Intelligent Tutoring Systems (ITS) that dynamically adapt to students, encouraging learning-focused behaviors and discouraging counterproductive ones. Furthermore, a deeper understanding of student behavior can lead to more robust student models, which better estimate knowledge states, misconceptions, and learning trajectories. For example, such a model could potentially detect when a student is overly reliant on LLM responses and flag this behavior to instructors in real-time, or be used to help develop assessments which are specifically designed to test for this over-reliance.
@@
In this work, we conduct a year-long case study to explore student behavior with an LLM, specifically in the context of upper-division, university-level Computer Science (CS) education. We develop a web application with the same core features and underlying LLM as ChatGPT, and deploy it in an introductory-level course on Artificial Intelligence (AI). We encourage students to use this tool, without restrictions, on programming assignments over the course of the semester. Across two semesters and with student consent we recorded 2,214 student-LLM conversations from 203 students across 7 unique assignments, totaling 16,851 unique utterances. We also collect 924 programming assignment submissions from 158 students who further agreed to share their work. In addition, we annotate these conversation utterances with a dialogue act (DA) labeling schema; the schema is inspired by existing studies on co-programming behavior and our systematic review of the conversations. We leverage an LLM prompting-based approach to do this annotation at scale and validate this approach with a human evaluation.
We explore how students engage with an LLM during programming assignments. In doing so our primary research objective is to uncover patterns of behavior across students and analyze if these trends correlate with course outcomes. To do this, we analyze these DA labels by performing regression and clustering analysis. Our key findings are: 1) students primarily use the LLM to ask questions, but our findings suggest that some question types are negatively correlated with outcome, suggesting student confusion possibly due to LLM hallucination; 2) certain behavior patterns such as asking conceptual questions or help in code writing are positively associated with both assignment and held out exam scores; 3) students with low and high usage of LLMs have similar average outcome but students that use that LLMs with high frequency tend to have less variance in a more narrow positive range; 4) clustering analysis reveals students who use the LLM primarily for code writing and not other writing requests tend to score higher on both assignments and exams. These findings provide data-driven insights into patterns of student LLM behaviors in a university-level CS course. We release this dataset to the broader research community to foster collaboration and exploration into LLMs’ impact on AI and CS education.
@@ 2.1 ChatGPT Classroom Studies
Closely related to this work are a handful of cases studies exploring student use of ChatGPT in university level CS classrooms. Ghimire et. al explore student use of ChatGPT for two assignments in a CS1 programming course, finding students often ask questions of the system instead of directly seeking solutions (Ghimire and Edwards, 2024). Ma et. al also perform a case study where CS1 students were allowed to use ChatGPT freely on in-class exercises and on a course project, but only for conceptual questions on homework citepma2024enhancing. Additionally, Brender et. al perform a case study on how graduate students in a robotics course utilize ChatGPT on a voluntary assignment, controlling for learning outcomes and clustering behavior into broad categories (Brender et al., 2024). Analysis of student prompts in these studies have noted a student trend to use models to ask questions more than directly writing code solutions which is consistent with our observations and those by Bhalerao (2024). Surveys on student perceptions towards LLMs have also been conducted which suggest that CS student adoption of LLMs is nuanced and not universal (Weber et al., 2024). Similarly, we observe that a some students use the system heavily while most students engage moderately with the model, revealing a range of adoption behaviors.
@@ 2.2 Education Discourse Analysis
It has been demonstrated by many that tutoring is a highly effective form of instruction Cohen et al. (1982). Thus, human tutoring dialogue has been studied extensively Graesser et al. (1995). In recent years, DA classification on conversation turns is often used in this discourse analysis Chen et al. (2011). DA labels classify a participant’s utterance by the function it serves in the conversation. When DA labeling is applied to an entire corpus of dialogues, it enables discourse-level analysis of conversation patterns that are frequent or significant. In education, significant patterns are often those associated with better learning outcomes, as such existing works have studied such patterns in tutor-student or student-student learning environments Graesser et al. (1995); Earle-Randell et al. (2023). Existing works also analyze human-machine classroom conversations between students and ITS chatbots Vail and Boyer (2014); Graesser et al. (2004); Fisher et al. (2020). These works provide promising insights which can inform more effective tutoring pedagogy Rus et al. (2017) and interactive learning tools Chen et al. (2011).
@@
While DA analysis is an attractive approach to discourse analysis, the labeling process poses many challenges, solutions to which constitute an large body of work Boyer et al. (2011); Tan et al. (2023); Ezen-Can and Boyer (2015); Lin et al. (2023); Boyer et al. (2010); Kumaran et al. (2023). First, there exists the task of annotating a large corpus of conversations. To automate such a task, various machine learning based approaches to this have been used such as Hidden Markov Models Boyer et al. (2010), decision trees Samei et al. (2014), and Markov Random Fields Ezen-Can and Boyer (2015). In our work, we employ a LLM prompt-based approach to DA labeling, similar to that of Wang et al. (2023). Second, there exists the challenge of robust schema design, where existing work has looked into refining prior labeling schemata to be less ambiguous Vail and Boyer (2014) and to try to develop classroom agnostic labeling schema Hennessy et al. (2016).
2.3 Education Dialogue Datasets
@@
In an effort to collectively improve data-driven learning technologies, there have been recent efforts to expand the amount of education dialogue datasets in a variety of learning domains. Datasets such as TalkMoves and NCTE directly capture and annotate classroom discourse in K-12 lessons Suresh et al. (2022); Demszky and Hill (2022). Other works employ crowd workers or leverage LLMs to simulate the student or teacher role, such as MathDial Macina et al. (2023), Teacher-Student Chatroom Corpus (TSCC) Caines et al. (2020), Conversational Instruction with Multi-responses and Actions (CIMA) Stasaski et al. (2020), and GrounDialog Zhang et al. (2023). Our work contributes to this growing body of research by providing a large-scale, annotated dialogue dataset of in a university-level CS course, offering insights into how students engage with LLM-chatbot tools in a real learning environment.
@@ 3 Dataset
Conceptual Questions Programming Language, Python Library, Computer Science, Programming Tools, Mathematics, Other Concept The user asks a general question about programming, libraries, tools, or CS concepts unrelated to the conversation.
@@
In this section, we detail the process by which the StudyChat dataset was formed and the classroom context.
@@ 3.1 Participants and Course Context
Our IRB-approved study was conducted among students aged 18 or above, who were enrolled in an AI course at a US R1 university in the Fall 2024 and Spring 2025 semesters. Across both semesters, 295 students were enrolled in the course, of which 203 agreed to participate in the study. Students who did not participate were also given equivalent access to use the tool over the semester, but their data was not collected. Students, were in their 3rd/4th years, had completed their foundational Computer Science course sequences and were taking higher-level elective courses.
Each week, students attend two 75-minute lectures and independently complete assignments related to the topics covered in lecture. The course consists of 3 modules taught over 15 weeks. The first module covers the foundations of the rational agent approach (focused on applied search algorithms), the second module covers fundamentals of machine learning with a focus on deep learning and neural networks, and the final module covers recent topics in applied AI. During each module, students complete 2-3 Python programming assignments related to the lecture material. In addition to the assignments, students were assessed via 2 end-of-module exams, and a final exam.
The programming assignments are Python-based and can be broadly categorized into two types. The first assignment type is a tutorial, where students are given incomplete Jupyter notebooks which guide students to learn fundamentals operations in applied machine learning Python libraries including Pandas, SciKit Learn, and PyTorch. The second assignment type is an applied scenario, where students are given Python starter code which they extend by applying an algorithm or concept taught in class. For example, in assignment 2, students are tasked with creating a text autocomplete engine by implementing various search algorithms to traverse graphs to find word completions. Assignments were altered slightly between semesters, but generally cover the same topics. All assignment instruction details, starter code, and 924 graded submissions are included in the dataset.
@@ 3.2 Data Collection
In order to effectively collect student utterances, we developed and deployed a web application designed to function similar to consumer LLM chatbot interfaces, shown in Figure 1. As our backing chatbot LLM, we used gpt-4o-mini, the LLM for the free tier of ChatGPT at the time of this study gpt . We evaluated this model by testing its capabilities with the libraries and concepts covered in the assignments, and found the model generally provides accurate guidance on the assignments and implementations that are often accurate. Students were instructed they could use this ChatGPT-based tool, without restrictions, on all assignments in the course. It was emphasized that their approach to using the tool would not affect their grade and they were encouraged to use the tool in any manner of their choosing. We used the default "You are a helpful assistant" system prompt, which does not discouraging answer-seeking behavior, and did not provide the chatbot with any course materials for reference. Our goal was to encourage students to use our platform instead of seeking an outside application for assignment help, and that the collected interactions would more accurately reflect how students generally use non-instructional, consumer chatbots.
@@ 3.3 Filtering of Student Identifiable Information
To remove any Personally Identifiable Information (PII) from the conversations and assignment submissions, we systematically analyze the conversations to uncover patterns of PII exposure. We first randomly sample 12 conversations from each course topic and read through the conversations, scanning for traces of PII. Our main finding of potentially sensitive student information is in the form of pasted terminal outputs and error messages, which sometimes contain PII, such as a user’s directory or GitHub account name. Compiling these common patterns, we then develop a regular expression-based script to remove these patterns and any direct matches to known user PII in the course roster. Across the entire dataset, this process removes 6,413 cases of possible PII exposure. Then, we sample 12 additional conversations from each topic and performed the same procedure. In our second pass of the second set of conversations, we did not find any additional patterns or cases of PII exposure. We acknowledge that although seemingly effective, this process does not guarantee the removal of all PII exposures.
@@ 3.4 Data Statistics
Across all conversations, we collect 16,851 student utterances and corresponding LLM responses over 7 graded assignments. The mean conversation length in the dataset is 7.6 utterances, with a median of 4.0 utterances per conversation. On average, each student has 83 utterances and 10.9 distinct conversations with the chatbot.
Table 2: Total counts per broad dialogue act label in StudyChat. Further breakdown of specific label counts can be found in Table 3.
Label f24 s25 Total
Conceptual Questions 2204 2995 5199
Writing Request 1664 2672 4336
Provide Context 1191 1837 3028
Contextual Questions 1216 1722 2938
Verification 293 354 647
Editing 153 265 418
Off Topic 128 122 250
Misc 15 20 35
@@ 3.5 Dialogue Act Annotation
In order to better understand common behavior and discourse-level trends in the conversations, we develop a DA labeling schema for student turns of the conversations. The schema is two-leveled, with 8 the broad labels and 31 specific labels, and is developed on top of existing work Wood et al. (2018) which conducted a “Wizard of Oz” experiment with programmers who were debugging alongside a simulated virtual assistant. It also mirrors our own observations after systematically reviewing the conversations, and aligns with those of Ma. et al Ma et al. (2024). The schema distinguishes the various ways students use an LLM to assist in writing aspects of their assignment and questions. We report total counts of each broad label in Table 2, total counts of specific labels in Table 3, and detail full descriptions of the schema in Table 1.
@@
To validate the accuracy of this labeling, we first sample a set of 150 utterances and develop an initial schema with labeling guidelines, including examples and descriptions for each DA. We then provide them to a research assistant, who independently labels the same set given our initial schema. We iterate on our labeling instructions and schema by computing the agreement between the two annotators and identifying sources of disagreement. After this iteration, our inter-rater agreement reached a Cohen’s kappa Cohen (1960) of 0.910 for broad level labels and 0.788 for specific labels, indicating almost perfect agreement on the broad level of DA and substantial agreement on the specific level Landis and Koch (1977). To validate this labeling, we provided these final instructions to two additional annotators who independently label the same set of utterances. We observed and average 4-way inter-rater agreement Cohen’s kappa of 0.740 for broad level agreement and 0.575 for specific level agreement, indicating substantial agreement on broad level labeling and moderate agreement on specific labeling.
@@
To apply this labeling at scale to the reminder of the dataset and since manual labeling is extraneous, we perform an LLM-prompting based approach similar to prior work Wang et al. (2023). We prompt the LLM with detailed instructions on the task, the annotation instructions, assignment instructions and the entire conversation up until the student turn to be labeled, and ask the LLM to provide the label based on our schema, along with a justification. We utilize GPT4.1 to perform this labeling, which we applied to the entire PII-filtered dataset.
@@
To validate the LLM-generated labels, we again systematically sample 154 utterances, distinct from those used in developing the initial labeling schema. We provide them to the same 4 human annotators without any further discussion or instructions. The resulting human inter-annotator agreement on the held out set is 0.578 on broad labels and 0.538 on specific labels, indicating moderate agreement between human annotators on the held out set. The average human-LLM agreement on the held out set is 0.5819 on the broad labels and 0.4927 on the specific labels, indicating moderate agreement on par with human inter-annotator agreement This drop in agreement is somewhat expected, since the new utterances in the held out set present new ambiguous, overlapping utterances to which multiple labels may apply.